The line-width of a bright fringe is sometimes defined as the separation between the points on the two sides of the central line where the intensity falls to half the maximum. Find the line-width of a bright fringe in a Young’s double slit experiment in terms of λ, d and D where the symbols have their usual meanings.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. i) When intensity is half the maximum l/l base max = 1/2 ⇒ 4a^2 cos^2(∅/2)/4a^2 = 1/4 ⇒ cos^2 (∅/2) = 1/4 ⇒ cos(∅/2) = 1/2 ⇒ ∅/2 = π/3 ⇒ ∅ = 2π/3 ⇒ Path difference, * = λ/3 ⇒ y = xD/d = λD/3d